What is the slope of a line that contains the point (2,-3) and has the same x-intercept as x+5y=8?
Answer by nerdybill(7384)

What is the slope of a line that contains the point (2,-3) and has the same x-intercept as x+5y=8?
.
First, find x-intercept of x+5y=8
Do this by setting y=0 and solve for y:
x+5y=8
0+5y=8
y = 8/5
So, the x-intercept is at (0, 8/5)
.
Now, we determine the slope between two points:
(2,-3) and (0, 8/5)
.
slope = (y2-y1)/(x2-x1)
slope = (8/5-(-3))/(0-2)
slope = (8/5+3)/(-2)
slope = (8/5+15/5)/(-2)
slope = (23/5)/(-2)
slope = -23/(5*2)
slope = -23/10
slope = -2.3